OpenStudy (samsan9):

which of the following equations represents a degenerate ellipse? A.) 9x^2+4y^2=1 B.) x^2/14+y^2/19=2 C.) x^2/10=-y^2/18 D.) all of the above are general ellipses

OpenStudy (samsan9):

i know it has something to do with zero and the center point but i never actually had to look for one before :/

OpenStudy (samsan9):

\[B.) \frac{ x^2 }{ 14 }+\frac{ y^2 }{ 19 }=2\]

OpenStudy (samsan9):

\[\frac{ x^2 }{ 10 }=-\frac{ y^2 }{ 18 }\]